# Lie algebras generated by reflections in types BCD

### v1

@@ 5.1. Statement of the main theorem
and hence $2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{\left\{\lambda,\pm\right\}})>\dim(S^{\left\{\lambda,\pm\right\}})-n$. As in (5.3.2) one sees that $\dim(S^{\left\{\lambda,\pm\right\}})\geq 2n$, and hence $02\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{\left\{\lambda,\pm\right\}})>$ 0. This implies first that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}$ is simple by [13, Lemme 15], and then that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}\cong\mathfrak{sl}\{\lambda,\pm\}$ by [13, Lemme 13]. ∎
@@
Let $n\geq 6$ be fixed, suppose that Theorem 5.1.1 is true for the value $n-1$.
@@
see [9, IX.5]. If $\rho_{J}\circ\rho_{I}^{-1}$ is of the first form, it follows that $S^{I}\cong S^{J}$ as $\mathfrak{b}_{n}^{\prime}$-modules, implying by Proposition 4.2.5 and the range of possible values for $I$ and $J$ in (1) that $I=J$. If $\rho_{J}\circ\rho_{I}^{-1}$ is of the second form, it follows that $S^{I}\cong[S^{J}]^{*,\operatorname{Lie}}$ as $\mathfrak{b}_{n}^{\prime}$-modules. However, since $[S^{(\lambda,\mu)}]^{*,\operatorname{Lie}}\cong S^{(\mu^{*},\lambda^{*})}$, and since the natural representation of $\mathfrak{sl}(k)$ is not self-dual for $k\geq 3$, it follows from Proposition 4.2.5 that $S^{I}\not\cong[S^{J}]^{*,\operatorname{Lie}}$ for any $I$ and $J$ in the range of possible values in (1). Thus the automorphism must be of the first form, and $I=J$.
@@
Let $n\geq 6$ be fixed, suppose that Theorem 5.1.1 is true for the value $n-1$. Then the (injective) maps (4.3.16) and (4.3.17) are surjections.

### v2

@@ 5.1. Statement of the main theorem
and hence $2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{\left\{\lambda,\pm\right\}})>\dim(S^{\left\{\lambda,\pm\right\}})-n$. As in (5.3.2) one sees that $\dim(S^{\left\{\lambda,\pm\right\}})\geq 2n$, and hence $2\cdot\operatorname{rk}(\mathfrak{h})-\dim(S^{\left\{\lambda,\pm\right\}})>0$. This implies first that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}$ is simple by [Marin:2007, Lemme 15], and then that $\mathfrak{d}_{\left\{\lambda,\pm\right\}}\cong\mathfrak{sl}\{\lambda,\pm\}$ by [Marin:2007, Lemme 13]. ∎
@@
Let $n\geq 6$ be fixed, suppose that Theorem 5.1.1 is true for the value n1n-1.
@@
see [Jacobson:1979, IX.5]. If $\rho_{J}\circ\rho_{I}^{-1}$ is of the first form, it follows that $S^{I}\cong S^{J}$ as $\mathfrak{b}_{n}^{\prime}$-modules, implying by Proposition 4.2.5 and the range of possible values for $I$ and $J$ in (1) that $I=J$. If $\rho_{J}\circ\rho_{I}^{-1}$ is of the second form, it follows that $S^{I}\cong[S^{J}]^{*,\operatorname{Lie}}$ as $\mathfrak{b}_{n}^{\prime}$-modules. However, since $S^{(\lambda,\mu)}]^{*,\operatorname{Lie}}\cong S^{(\mu^{*},\lambda^{*})}$, and since the natural representation of $\mathfrak{sl}(k)$ is not self-dual for $k\geq 3$, it follows from Proposition 4.2.5 that $S^{I}\not\cong[S^{J}]^{*,\operatorname{Lie}}$ for any $I$ and $J$ in the range of possible values in (1). Thus the automorphism must be of the first form, and $I=J$.
@@
Let $n\geq 6$ be fixed, suppose that Theorem 5.1.1 is true for the value n1n-1. Then the (injective) maps (4.3.16) and (4.3.17) are surjections.